OpenStudy (anonymous):

HEllllllP Please On the graph of the equation 2x + 5y = -30, what is the value of the y-intercept? What is the standard form equation of the line passing through (1, -5) with a slope of -1?

OpenStudy (anonymous):

y = -6

OpenStudy (anonymous):

here's another What is the standard form equation of the line passing through (1, -5) with a slope of -1?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you show the steps for the second one

OpenStudy (anonymous):

y - - 5 = -1(x-1) y+5 = -1x + 1 y= -x - 4 The answer is x + y = –4

OpenStudy (anonymous):

ohh, i see just put it in to pint slope form then convert to standard

OpenStudy (anonymous):

What is the equation of the line passing through (-3, 6) and parallel to the line y = 4x - 2 in standard form?

OpenStudy (anonymous):

hey i've got 2 more after this think you can help

OpenStudy (anonymous):

4x - y= -18

OpenStudy (anonymous):

And sure

OpenStudy (anonymous):

What is the equation of the line passing through (2, -2) and perpendicular to the line y = -x + 3 in slope-intercept form?

OpenStudy (anonymous):

Solve x - 2y = 4 for y.

OpenStudy (anonymous):

First one is y = x - 4

OpenStudy (anonymous):

y=(x-4)/2

OpenStudy (anonymous):

thats what i came up with but the possible answer are these a. y = x - 2 b. y = -x + 4 c. y = x - 2 d. y = x + 4

OpenStudy (anonymous):

a. y = x - 2 b. y = -x + 4 c. y = x - 2 d. y = x + 4

OpenStudy (anonymous):

oops one sec

OpenStudy (anonymous):

a. y = x - 2 b. y = -x + 4

OpenStudy (anonymous):

c. y=onehalfx-2

OpenStudy (anonymous):

Oh okay hold on

OpenStudy (anonymous):

y=-onehalfx+4

OpenStudy (anonymous):

forget the first abcd, the

OpenStudy (anonymous):

What are you answer choices for the question?

OpenStudy (anonymous):

& which question are you talking about?

OpenStudy (anonymous):

i am talking about solve x - 2y = 4 for y., and the answers are these a. y = x - 2 b. y = -x + 4 c.y=onehalfx-2 d.y=-onehalfx+4

OpenStudy (anonymous):

starbelly you got it

OpenStudy (anonymous):

Yeah, one sec

OpenStudy (anonymous):

y= -1/2x - 2

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

it was wrong, i jsut submitted and it was wrong

OpenStudy (anonymous):

Oh. I'm was just redoing the problem but sorry